# A signal separation view of classification

### v1

@@ Abstract
The problem of classification in machine learning has often been approached in terms of function approximation. In this paper, we propose an alternative approach for classification in arbitrary compact metric spaces which, in theory, yields both the number of classes, and a perfect classification using a minimal number of queried labels. Our approach uses localized trigonometric polynomial kernels initially developed for the point source signal separation problem in signal processing. Rather than point sources, we argue that the various classes come from different probability distributions. The localized kernel technique developed for separating point sources is then shown to separate the supports of these distributions. This is done in a hierarchical manner in our MASC algorithm to accommodate touching/overlapping class boundaries. We illustrate our theory on several simulated and real life datasets, including the Salinas and Indian Pines hyperspectral datasets and a document dataset.
@@ 1 Introduction
A fundamental problem in machine learning is the following. Let $\{(x_{j},y_{j})\}_{j=1}^{M}$ be random samples from an unknown probability distribution $\tau$. The problem is to approximate the conditional expectation $f(x)=\mathbb{E}_{\tau}(y|x)$ as a function of $x$. Naturally, there is a huge amount of literature studying function approximation by commonly used tools in machine learning such as neural and kernel based networks. For example, the universal approximation theorem gives conditions under which a neural network can approximate an arbitrary continuous function on an arbitrary compact subset of the ambient Euclidean space. The estimation of the complexity of the approximation process typically assumes some smoothness conditions on $f$, examples of which include, the number of derivatives, membership in various classes such as Besov spaces, Barron spaces, variation spaces, etc.
@@
A very important problem is one of classification. Here the values of $y_{j}$ can take only finitely many (say $K$) values, known as the class labels. In this case, it is fruitful to approximate the classification function, defined by $f(x)=\operatorname*{argmax}_{k}\mathsf{Prob}(k|x)$ [23]. Obviously, this function is only piecewise continuous, so that the universal approximation theorem does not apply directly. In the case when the classes are supported on well separated sets, one may refer to extension theorems such as Stein extension theorems [28] in order to justify the use of the various approximation theorems to this problem.
@@
While these arguments are sufficient for pure existence theorems, they also create difficulties in an actual implementation, in particular, because these extensions are not easy to construct. In fact, this would be impossible if the classes are not well separated, and might even overlap. Even if the classes are well separated, and each class represents a Euclidean domain, any lack of smoothness in the boundaries of these domains is a problem. Some recent efforts, for example, by Petersen and Voigtländer [24] deal with the question of accuracy in approximation when the class boundaries are not smooth. However, a popular assumption in the last twenty years or so is that the data is distributed according a probability measure supported on a low dimensional manifold of a high dimensional ambient Euclidean space. In this case, the classes have boundary of measure 0 with respect to the Lebesgue measure on the ambient space. Finally, approximation algorithms, especially with deep networks, utilize a great deal of labeled data.
@@
In this paper, we propose a different approach as advocated in [7]. Thus, we do not assume that $\mathsf{Prob}(k|x)$ is a function, but assume instead that the points $x_{j}$ in class $k$ comprise the support of a probability measure $\mu_{k}$. The marginal distribution $\mu$ of $\tau$ along $x$ is then a convex combination of the measures $\mu_{k}$. The fundamental idea is to determine the supports of the measures $\mu_{k}$ rather than approximating $\mu_{k}$’s themselvesiiiIf $\nu$ is a positive measure on a metric space $\mathbb{M}$, we define the support of any positive measure $\nu$ by $\mathsf{supp}(\nu)=\{x\in\mathbb{M},\nu({\mathbb{B}}(x,r)>0\mbox{ for all }r>0\}$, where ${\mathbb{B}}(x,r)$ is the ball of radius $r$ centered at $x$.. This is done in an unsupervised manner, based only on the $x_{j}$’s with no label information. Having done so, we may then query an oracle for the label for one point in the support of each measure, which is then necessarily the label for every other point in the support. Thus, we aim to achieve in theory a perfect classification using a minimal amount of judiciously chosen labeled data.
@@
In order to address the problem of overlapping classes, we take a hierarchical multiscale approach motivated by a paper [5] of Chaudhury and Dasgupta. Thus, for each value $\eta$ of the minimal separation among classes, we assume that the support of $\mu$ is a disjoint union of $K_{\eta}$ subsets, each representing one of $K_{\eta}$ classes, leaving an extra set, representing the overlapping region. When we decrease $\eta$, we may eventually capture all the classes, leaving only a negligible overlapping region (ideally with $\mu$-probability 0).
@@
In [7], we had the epiphany that the problem is analogous to the problem of point source signal separation. If each $\mu_{k}$ were a Dirac delta measure supported at say $\omega_{k}$, the point source signal separation problem is to find these point sources from finitely many observations of the Fourier transform of $\mu$. In the classification problem we do not have point sources and the information comprises samples from $\mu$ rather than its Fourier transform. Nevertheless, we observed in [7] that the techniques developed for the point source signal separation problem can be adapted to the classification problem viewed as the the problem of separation of the supports of $\mu_{k}$. In that paper, we assumed only that the data is supported on a compact subset of a Eulidean space, and used a specially designed localized kernel based on Hermite polynomials [6] for this purpose. Since Hermite polynomials are intrinsically defined on the whole Euclidean space, this creates both numerical and theoretical difficulties. In this paper, we allow the data to come from an arbitrary compact metric space, and use localized trigonometric polynomial kernels instead. We feel that this leads to a more satisfactory theory, although one of the accomplishments of this paper is to resolve the technical difficulties required to achieve this generalization.
@@
Our work belongs in the general theory of active learning. In Section 2, we review some literature in this area which is somewhat related to the present work. In Section 3, we give a brief discussion of the point source signal separation problem and the use of localized trigonometric polynomial kernels to solve it. In Section 4, we describe the background needed to formulate our theorems, which are given in Section 5. The algorithm MASC to implement these results in practice is given in Section 6, and illustrated in the context of a simulated circle and ellipse data set, a document dataset, and two hyperspectral datasets. The proofs of the results in Section 5 are given in Section 8.
@@ 2 Related works
Perhaps the most relevant work to this paper is that of [7]. That paper also outlines the theory and an algorithm for a classification procedure using aa thresholding set based on a localized kernel. There are three major improvements we have made relative to that work in this paper:
@@
    We have constructed the kernel in this paper in terms of trigonometric functions, whereas in [7] the kernel was constructed from Hermite polynomials. The trigonometric kernel is much faster in implementations for two reasons: 1) each individual polynomial is extremely quick to compute and 2) the trigonometric kernel deals only with trigonometric polynomials up to degree $n$, whereas the Hermite polynomial based kernel needs polynomials up to degree $n^{2}$ to achieve the same support estimation bounds.
@@
    This paper deals with arbitrary compact metric spaces (allowing for a rescaling of the data so that the maximum distance between values is $\leq\pi$), whereas [7] dealt with compact subsets of the Euclidean space and had a requirement on the degree of the kernel dependent upon the diameter of the data in terms of Euclidean distance.
@@
    In [7], an algorithm known as Cautious Active Clustering (CAC) was developed. In this paper we present a new algorithm with several implementation advantages over CAC. We discuss this topic in more depth in Section 6.2.
Another pair of related works is that of the Learning by Active Nonlinear Diffusion (LAND) and Learning by Evolving Nonlinear Diffusion (LEND) algorithms [15, 31]. Like the present work, these algorithms use a kernel-based density estimation when deciding points to query. However, LAND and LEND both use a Gaussian kernel applied on $k$ neighbors for the density estimation and weight it by a diffusion value. Then, the queried points are simply those with the highest of the combined weights. The diffusion value corresponds to a minimal diffusion distance among points with a higher density estimation. For the point with the maximal density estimation, a maximal diffusion distance among other data points is taken as the weight. This extra weighting procedure is absent from our theory and algorithm, which uses an estimation approach based purely on a localized kernel to decide on points to query. In our algorithm, we take a multiscale approach and decide on query points at each level instead of a global listing of the data points.
In [32], an active learning approach using neural networks is developed. This work focuses on binary classification and developing models using a neural network framework such that a sufficient number of queries will achieve a desired accuracy.
A study of two types of uncertainty in active learning problems is discussed in [27]. The two critical types of uncertainty are 1) a data point is likely to belong to multiple labels, 2) a data point is not likely to belong to any label. Our work also seeks to distinguish between data points which are uncertain in the second sense, using a graph construction approach and potentially also a thresholding set for high-density points. When our algorithm encounters points which are uncertain in the first sense, it elects not to assign a label right away, instead coming back to it once the “confident” points have been classified.
@@
Our method and algorithm is meant to work on general data sampled from compact metric spaces. One difficulty that algorithms may face is the presence of highly imbalanced data (i.e. where some class labels dominate over others in a data set). The problem of tackling this difficulty is studied in [29], where an approach to querying imbalanced data using a balance of two principles is employed: exploration and exploitation. During the exploration phase, the algorithm seeks out points to query in low-sought regions. During the exploitation phase, the algorithm seeks to query points in the most critical explored regions. Our algorithm works in a different fashion, by querying points which we believe to be in high-density portions of a label’s support and extending the label to nearby points until it “bumps” against points which may belong to another label.
We list the survey by Tharwat and Schenck [30] as a resourceful survey of recent developments in active learning.
@@ 3 Point source signal separation
The problem of signal separation goes back to early work of de Prony [8], and can be stated as: estimate the coefficients $a_{k}$ and locations $\omega_{k}$ constituting $\mu=\sum_{k=1}^{K}a_{k}\delta_{\omega_{k}}$, from observations of the form
@@
There is much literature on methods to approach this problem, and we cite [25] as a text one can use to familiarize themselves with the topic. If we assume $\omega_{k}=k\Delta$ for some $\Delta\in\mathbb{R}^{+}$ and allow measurements for any $x\in[-\Omega,\Omega]$ for some $\Omega\in\mathbb{R}^{+}$, then recovery is possible so long as we are above the Rayleigh threshold, i.e. $\Omega\geq\pi/\Delta$ [9]. The case where this threshold is not satisfied is known as super-resolution. Much further research has gone on to investigate the super-resolution problem, such as [1, 4, 14].
We now introduce a particular method of interest for signal separation from [22] and further developed in [18]. The method takes the following approach to estimate the coefficients and locations of $\mu$, without the assumption that the $\omega_{k}$’s should be at grid points, and the additional restriction that only finitely many integer values of $x$ are allowed. We start with the trigonometric moments of $\mu$:
@@
A key property of $\Phi_{n}$ is the localization property (cf. [10, 13], where the notation is different): For any integer $S\geq 3$,
@@
The theoretical details of this sentiment are described more rigorously in [10, 13]. Here, we only give two examples to illustrate.
@@
In Figure 1, we show the graphs of the “power spectrum” $|\sigma_{n}(\mu)(x)|$ for $n=64,256$.
Figure 1: Left: $|\sigma_{256}(\mu)(x)|$ has peaks at the points 1,2,2.05-1,2,2.05, and is small everywhere else. Vertical red lines indicate the positions of these points. Middle: A close-up view of $|\sigma_{256}(\mu)(x)|$ near $x=2$ to show an accurate detection of the close-by points $2,2.05$. Right: A close-up view of $|\sigma_{64}(\mu)(x)|$ near $x=2$ to show the non-detection of the close-by points $2,2.05$.
We see from the leftmost figure that $\left|\sigma_{256}\right|$ has peaks at approximately 1,2,2.05-1,2,2.05, and is very small everywhere else on $\pi,\pi]$. The middle figure is a close-up view to highlight an accurate detection of the close-by point sources $2,2.05$. The rightmost figure shows that with $n=64$, such a resolution is not possible. When we wish to automate this, we need to figure out a threshold so that we should look only at peaks above the threshold. As the middle figure shows, there are sidelobes around each peak (and in fact, small sidelobes at many other places on $\pi,\pi]$). In theory, this threshold is $\min_{k}|a_{k}|/2$, which we do not know in practice. If we set it too low, then we might “detect” non-existent point sources near $2,2.05$. On the other hand, if we set it too high, then we would lose the low amplitude point source at 1-1. Some ideas on how to set an appropriate threshold, especially in the presence of noise are discussed in [13]. Another important quantity is the minimal separation $\eta=\min_{k\neq j}|(\omega_{k}-\omega_{j})\mbox{ mod }2\pi|$ among the point sources. As the middle and right figures show, the detection of point sources which are very close-by requires the knowledge of a larger number of moments. It is shown in [16] that one must have $n\gtrsim\eta^{-1}$ in order to have sufficient resolution to recover the point sources in a stable manner. ∎
@@
We define a distribution $\mu$ on $\mathbb{T}$} as a convex combination of:
@@
  • Three point-mass distributions at 2,0.4,1.5-2,0.4,1.5, with weights of $60/3900,120/3900,120/3900$ respectively (anomaly).
We take 3900 samples from this distribution (the number of points from each part of the distribution corresponding to the numerator of the weight). The samples from the distribution are visualized in Figure 2 (a) as a normalized histogram. Then, we apply $\sigma_{128}$ to get an estimation of the support, as seen in Figure 2 (b). Not only do we get an idea of the support of the distribution by looking at $\sigma_{128}$, but also the amplitudes of the non-atomic components of the distribution. Since we are dealing with finitely many samples, we in fact are only estimating the integral in the definition of $\sigma_{n}$ as a Monte-Carlo type summation. That is, with data $\{u_{j}\}_{j=1}^{M}$ sampled randomly from $\mu$, we estimate

### v2

@@ Abstract
The problem of classification in machine learning has often been approached in terms of function approximation. In this paper, we propose an alternative approach for classification in arbitrary compact metric spaces which, in theory, yields both the number of classes, and a perfect classification using a minimal number of queried labels. Our approach uses localized trigonometric polynomial kernels initially developed for the point source signal separation problem in signal processing. Rather than point sources, we argue that the various classes come from different probability measures. The localized kernel technique developed for separating point sources is then shown to separate the supports of these measures. This is done in a hierarchical manner in our MASC algorithm to accommodate touching/overlapping class boundaries. We illustrate our theory on several simulated and real life datasets, including the Salinas and Indian Pines hyperspectral datasets and a document dataset.
@@ 1 Introduction
A fundamental problem in machine learning is the following. Let $\{(x_{j},y_{j})\}_{j=1}^{M}$ be random samples from an unknown probability measure $\tau$. The problem is to approximate the conditional expectation $f(x)=\mathbb{E}_{\tau}(y|x)$ as a function of $x$. Naturally, there is a huge amount of literature studying function approximation by commonly used tools in machine learning such as neural and kernel based networks. For example, the universal approximation theorem gives conditions under which a neural network can approximate an arbitrary continuous function on an arbitrary compact subset of the ambient Euclidean space [chuili1992, cybenko-net]. The estimation of the complexity of the approximation process typically assumes some smoothness conditions on $f$, examples of which include, the number of derivatives, membership in various classes such as Besov spaces, Barron spaces, variation spaces, etc [barron-universal, mhaskar-multilayer, mhaskar-net].
@@
A very important problem is one of classification. Here the values of $y_{j}$ can take only finitely many (say $K$) values, known as the class labels. In this case, it is fruitful to approximate the classification function, defined by $f(x)=\operatorname*{argmax}_{k}\mathsf{Prob}(k|x)$ [murphy2022probabilistic]. Obviously, this function is only piecewise continuous, so that the universal approximation theorem does not apply directly. In the case when the classes are supported on well-separated sets, one may refer to extension theorems such as Stein extension theorems [stein] in order to justify the use of the various approximation theorems to this problem.
@@
While these arguments are sufficient for pure existence theorems, they also create difficulties in an actual implementation, in particular, because these extensions are not easy to construct. In fact, this would be impossible if the classes are not well-separated, and might even overlap. Even if the classes are well-separated, and each class represents a Euclidean domain, any lack of smoothness in the boundaries of these domains is a problem. Some recent efforts, for example, by Petersen and Voigtländer [petersen-relu] deal with the question of accuracy in approximation when the class boundaries are not smooth. However, a popular assumption in the last twenty years or so is that the data is distributed according a probability measure supported on a low dimensional manifold of a high dimensional ambient Euclidean space. In this case, the classes have boundary of measure 0 with respect to the Lebesgue measure on the ambient space. Finally, approximation algorithms, especially with deep networks, utilize a great deal of labeled data.
@@
In this paper, we propose a different approach as advocated in [cloningercluster]. Thus, we do not assume that $\mathsf{Prob}(k|x)$ is a function, but assume instead that the points $x_{j}$ in class $k$ comprise the support of a probability measure $\mu_{k}$. The marginal distribution $\mu$ of $\tau$ along $x$ is then a convex combination of the measures $\mu_{k}$. The fundamental idea is to determine the supports of the measures $\mu_{k}$ rather than approximating $\mu_{k}$’s themselvesiiiIf $\nu$ is a positive measure on a metric space $\mathbb{M}$, we define the support of any positive measure $\nu$ by $\mathsf{supp}(\nu)=\{x\in\mathbb{M},\nu({\mathbb{B}}(x,r))>0\mbox{ for all }r>0\}$, where ${\mathbb{B}}(x,r)$ is the ball of radius $r$ centered at $x$.. This is done in an unsupervised manner, based only on the $x_{j}$’s with no label information. Having done so, we may then query an oracle for the label of one point in the support of each measure, which is then necessarily the label for every other point in the support. Thus, we aim to achieve in theory a perfect classification using a minimal amount of judiciously chosen labeled data.
@@
In order to address the problem of overlapping classes, we take a hierarchical multiscale approach motivated by a paper [dasgupta2010] of Chaudhury and Dasgupta. Thus, for each value $\eta$ of the minimal separation among classes, we assume that the support of $\mu$ is a disjoint union of $K_{\eta}$ subsets, each representing one of $K_{\eta}$ classes, leaving an extra set, representing the overlapping region. When we decrease $\eta$, we may eventually capture all the classes, leaving only a negligible overlapping region (ideally with $\mu$-probability 0).
@@
In [cloningercluster], we explored a new insight that the problem is analogous to the problem of point source signal separation. If each $\mu_{k}$ were a Dirac delta measure supported at say $\omega_{k}$, the point source signal separation problem is to find these point sources from finitely many observations of the Fourier transform of $\mu$. In the classification problem we do not have point sources and the information comprises samples from $\mu$ rather than its Fourier transform. Nevertheless, we observed in [cloningercluster] that the techniques developed for the point source signal separation problem can be adapted to the classification problem viewed as the the problem of separation of the supports of $\mu_{k}$. In that paper, we assumed only that the data is supported on a compact subset of a Eulidean space, and used a specially designed localized kernel based on Hermite polynomials [chuigaussian] for this purpose. Since Hermite polynomials are intrinsically defined on the whole Euclidean space, this creates both numerical and theoretical difficulties. In this paper, we allow the data to come from an arbitrary compact metric space, and use localized trigonometric polynomial kernels instead. We feel that this leads to a more satisfactory theory, although one of the accomplishments of this paper is to resolve the technical difficulties required to achieve this generalization.
Our work belongs to the general theory of active learning. In the active learning paradigm for machine learning, one is only given $\mathcal{D}\coloneqq\{x_{j}\}_{j=1}^{M}$. However, for any $x_{j}$ one is allowed to query an oracle for the true value $y_{j}$ associated with it. The understanding is that this process is costly, so one only wants to query as few points as possible to accurately attain the $y$ values for the rest of the data set. In this way, active learning exists as a bridge between unsupervised learning (where no $y_{j}$ values are available) and semi-supervised learning (where the $y_{j}$ values are available only on a preselected set of points $x_{j}$). The critical problem of active learning is to decide on which points to query. One wishes to query points that give as much information on the rest of the data as possible. We list the survey by Tharwat and Schenck [tharwat-survey] as a resource for recent developments in active learning.
Our main theorem 5.2 concerns a method to estimate, based on the finite data $\mathcal{D}$, the supports of a measure corresponding to different class labels. To aid in this estimation, we assume that the data lies on an unknown subset of a known compact metric space $\mathbb{M}$ with metric $\rho$. The idea of the theorem is shown as a simple visual in Figure 1. In this figure, we see that the true classes (shown at the top) have no minimal separation. Supposing we can define partitions $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta},\mathbf{S}_{3,\eta}$ where $\mathbf{S}_{3,\eta}$ is small ($\mu(\mathbf{S}_{3,\eta})\to 0$ as $\eta\to 0$) and $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta}$ correspond to the original class labels and have separation $2\eta$, then our theorem gives conditions (based on some parameters $n,\Theta$ and size of the set $\mathcal{D}$) for when our support estimation sets $\mathcal{G}_{1,\eta,n}(\Theta),\mathcal{G}_{2,\eta,n}(\Theta)$ have separation $\eta$ and closely estimate $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta}$. In this paper, we consider measures which satisfy this partitioning property for any sufficiently small $\eta$. This allows us to take $\eta\to 0$ so that $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta}$ approach the true classes. This framework allows us to consider the machine learning classification problem even for classes that may have no minimal separation.
Figure 1: Visualization of our main theorem. Top: Supports of two classes with no minimal separation. Bottom: The two classes are separated into sets $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta}$ (blue and green) with separation $2\eta$ by removing a remainder set $\mathbf{S}_{3,\eta}$ (red). Our theorem gives conditions for when our support estimation sets $\mathcal{G}_{1,\eta,n}(\Theta),\mathcal{G}_{1,\eta,n}$ (light blue and light green) have separation $\eta$ and are close estimations of $\mathbf{S}_{1,\eta},\mathbf{S}_{2,\eta}$ respectively.
Our Algorithm 1, which has computational improvements and demonstrated accuracy improvements from our prior work in [cloningercluster], uses the theorem as a starting point to classify data sets in a multiscale active learning fashion. We start by thresholding away low-density points using a measure support estimator funtion (defined in (5.1)). Then, we iteratively create successively larger discconnected graphs among the high-density points to extend queried labels in a cautious manner. By “cautious”, we mean that the process is halted locally in the event of conflicting labels belonging to a single graph component. For any graph component where we have not yet queried a point for a label, we do so by choosing the point in the component which maximizes the same support estimator function.
@@
In Section 2, we review some literature in this area which is somewhat related to the present work. In Section 3, we give a brief discussion of the point source signal separation problem and the use of localized trigonometric polynomial kernels to solve it. In Section 4, we describe the background needed to formulate our theorems, which are given in Section 5. The algorithm MASC to implement these results in practice is given in Section 6, and illustrated in the context of a simulated circle and ellipse data set, a document dataset, and two hyperspectral datasets. The proofs of the results in Section 5 are given in Section 8.
@@ 2 Related works
Perhaps the most relevant work to this paper is that of [cloningercluster]. That paper also outlines the theory and an algorithm for a classification procedure using a thresholding set based on a localized kernel. There are three major improvements we have made relative to that work in this paper:
@@
    We have constructed the kernel in this paper in terms of trigonometric functions, whereas in [cloningercluster] the kernel was constructed from Hermite polynomials. The trigonometric kernel is much faster in implementations for two reasons: 1) each individual polynomial is extremely quick to compute and 2) the trigonometric kernel deals only with trigonometric polynomials up to degree $n$, whereas the Hermite polynomial based kernel needs polynomials up to degree $n^{2}$ to achieve the same support estimation bounds.
@@
    This paper deals with arbitrary compact metric spaces (allowing for a rescaling of the data so that the maximum distance between values is $\leq\pi$), whereas [cloningercluster] dealt with compact subsets of the Euclidean space and had a requirement on the degree of the kernel dependent upon the diameter of the data in terms of Euclidean distance.
@@
    In [cloningercluster], an algorithm known as Cautious Active Clustering (CAC) was developed. In this paper we present a new algorithm with several implementation advantages over CAC. We discuss this topic in more depth in Section 6.2.
An important aspect of active learning is how samples are queried to minimize uncertainty. A study of two types of uncertainty in active learning problems is discussed in [sharma-active]. The two critical types of uncertainty are 1) a data point is likely to belong to multiple labels, 2) a data point is not likely to belong to any label. In our algorithm, we deal with uncertain points after we have finished querying and extending the graph components. At this point, the only unlabeled points will be those that belonged to clusters with conflicting queried labels (uncertainty of the first type) or thresholded out at the beginning (uncertainty of the second type). We utilize all of the information from the iterative portion of our algorithm to then assign these points labels in a semi-supervised fashion.
@@
One difficulty that algorithms may face in the active learning setting is the presence of highly imbalanced data (i.e. where data associated with some class labels is much more plentiful than others). In [tharwat-active], the authors discuss two concepts that an active learning algorithm should employ to be successful: exploration and exploitation. During the exploration phase, their algorithm seeks out points to query in low-sought regions. During the exploitation phase, their algorithm seeks to query points in the most critical explored regions. Our algorithm balances these principles in a different way: by querying points which we believe to be in high-density portions of a label’s support (exploitation) and extending the label to nearby points until it “bumps” against points which may belong to another label (exploration).
There are many other approaches to active learning that we categorize broadly into three groups: diffusion geometry, graph-based approaches, and neural networks. Our work has significant overlap with diffusion geometry and graph-based approaches, but no direct ties to neural networks. We list [tharwat-survey] as a resourceful survey of recent developments in active learning.
Graph-based algorithms:
Graph-based active learning is a broad category of active learning algorithms which leverage graph structures for clustering. The broad approach for many of these algorithms is two-part: 1) to come up with some weighting between unlabeled data points based on the known labels to construct a weighted graph 2) minimizing or maximizing some function over the nodes of the graph to decide on the next point to query. The algorithms will switch between steps 1 and 2 iteratively to query more points and improve the classification. In [calder-laplace, chen-batch, miller-uncertainty, bertozzi-model-change], for instance, weight functions based on the graph Laplacian are applied to the data and minimized among the unlabeled data to choose successive query points. The weight function is chosen so as to decay away from the labeled data in a manner to quantify the uncertainty or information-gain among the unlabeled data points. In [miller-dirichlet], the function is chosen to be a Dirichlet variance with the purpose of representing Bayesian information regarding plausible labels for the unlabeled data. These approaches use successive querying schemes, where each new query point requires a new calculation across all of the unlabeled data.
Our method also uses graph clustering. In contrast to many algorithms in this setting though, we intentionally look at disconnected graphs and, in particular, extend queried labels to all of the points in each of the graph components. Since our work assumes a known metric $\rho$, the distance between points by this metric works as a similarity scheme for the purpose of graph construction. Furthermore, we use the graph structure only as a clustering tool, and do not decide on points to query via any weights defined on edges. Rather, we use a density estimator to query high density points in each of the clusters we find in a hierarchical manner.
Diffusion Geometry:
Diffusion geometry active learning can be considered as a subset of graph-based active learning, where the edge weights are decided by a diffusion distance, given in terms of a Markov transition matrix. Unlike the graph-based algorithms above which iteratively query points in a two-step process, the diffusion geometry methods mentioned here create an ordering of the data points based on the weights and do the querying based on this ordering all in one step. In Section 7, we have compared our algorithm with two state-of-the-art methods: Learning by Active Nonlinear Diffusion (LAND) and Learning by Evolving Nonlinear Diffusion (LEND) algorithms [murphy-land, murphy-lend]. We note that there also exist results from a diffusion geometry perspective focused on unsupervised learning [murphy-unsupervised, polk-multiscale], but both works show the potential improvements to be gained from an active learning framework. Like the present work, diffusion geometry approaches use a kernel-based density estimation when deciding points to query. However, LAND and LEND both use a Gaussian kernel applied on $k$ neighbors for the density estimation and weight it by a diffusion value. Then, the queried points are simply those with the highest of the combined weights. The diffusion value corresponds to a minimal diffusion distance among points with a higher density estimation. For the point with the maximal density estimation, a maximal diffusion distance among other data points is taken as the weight. This extra weighting procedure is absent from our theory and algorithm, which uses an estimation approach based purely on a localized kernel to decide on points to query. In our algorithm, we take a multiscale approach and decide on query points at each level instead of a global listing of the data points.
Neural Networks:
Neural network approaches broadly decide on points to query by inputting the data into a neural network. Although our approach is disparate from neural networks, we highlight a few recent advances in this area. In [nowak-network], an active learning approach using neural networks is developed. This work focuses on binary classification and developing models using a neural network framework such that a sufficient number of queries will achieve a desired accuracy. In [schroder-network], transformers are utilized in the active learning process, with particular emphasis on text document classification applications. In [kim-network], a method to effectively augment data sets with additional “labeled” data points is investigated for deep active learning.
@@ 3 Point source signal separation
The problem of signal separation goes back to early work of de Prony [prony], and can be stated as: estimate the coefficients $a_{k}$ and locations $\omega_{k}$ constituting $\mu=\sum_{k=1}^{K}a_{k}\delta_{\omega_{k}}$, from observations of the form
@@
There is much literature on methods to approach this problem, and we cite [plonkafourier] as a text one can use to familiarize themselves with the topic. If we assume $\omega_{k}=k\Delta$ for some $\Delta\in\mathbb{R}^{+}$ and allow measurements for any $x\in[-\Omega,\Omega]$ for some $\Omega\in\mathbb{R}^{+}$, then recovery is possible so long as we are above the Rayleigh threshold, i.e. $\Omega\geq\pi/\Delta$ [donoho-superres]. The case where this threshold is not satisfied is known as super-resolution. Much further research has gone on to investigate the super-resolution problem, such as [batenkov-superres, candes-superres, li-superres].
We now introduce a particular method of interest for signal separation from [loctrigwave] and further developed in [mhaskar-kitimoon-raj]. The method takes the following approach to estimate the coefficients and locations of $\mu$, without the assumption that the $\omega_{k}$’s should be at grid points, and the additional restriction that only finitely many integer values of $x$ are allowed. We start with the trigonometric moments of $\mu$:
@@
A key property of $\Phi_{n}$ is the localization property (cf. [mhaskar-prestin-filbir, kitimoon2025localized], where the notation is different): For any integer $S\geq 3$,
@@
The theoretical details of this sentiment are described more rigorously in [mhaskar-prestin-filbir, kitimoon2025localized]. Here, we only give two examples to illustrate.
@@
In Figure 2, we show the graphs of the “power spectrum” $|\sigma_{n}(\mu)(x)|$ for $n=64$ and $n=256$.
Figure 2: Left: $|\sigma_{256}(\mu)(x)|$ has peaks at the points 1,2,2.05-1,2,2.05, and is small everywhere else. Vertical red lines indicate the positions of these points. Right: A close-up view of $|\sigma_{256}(\mu)(x)|$ near $x=2$ to show an accurate detection of the close-by points $2,2.05$. Bottom: A close-up view of $|\sigma_{64}(\mu)(x)|$ near $x=2$ to show the non-detection of the close-by points $2,2.05$.
We see from the figure on the left that $\left|\sigma_{256}\right|$ has peaks at approximately 1,2,2.05-1,2,2.05, and is very small everywhere else on $\pi,\pi]$. The figure on the right is a close-up view to highlight an accurate detection of the close-by point sources $2,2.05$. The figure on the bottom shows that with $n=64$, such a resolution is not possible. When we wish to automate this, we need to figure out a threshold so that we should look only at peaks above the threshold. As the middle figure shows, there are sidelobes around each peak (and in fact, small sidelobes at many other places on $\pi,\pi]$). In theory, this threshold is $\min_{k}|a_{k}|/2$, which we do not know in practice. If we set it too low, then we might “detect” non-existent point sources near $2,2.05$. On the other hand, if we set it too high, then we would lose the low amplitude point source at 1-1. Some ideas on how to set an appropriate threshold, especially in the presence of noise are discussed in [kitimoon2025localized]. Another important quantity is the minimal separation $\eta=\min_{k\neq j}|(\omega_{k}-\omega_{j})\mbox{ mod }2\pi|$ among the point sources. As the middle and right figures show, the detection of point sources which are very close-by requires the knowledge of a larger number of moments. It is shown in [beyondsuper] that one must have $n\gtrsim\eta^{-1}$ in order to have sufficient resolution to recover the point sources in a stable manner. ∎
Remark 3.1.
We make some remarks about the notion of minimal separation introduced in Example 3.1. In the case of uniform sampling, the minimal separation takes on the same value as the sampling rate $\Delta$ from before. However, this perspective of minimal separation also allows one to consider the case of non-uniform sampling. The analogue of the Rayleigh limit is a theorem in [beyondsuper] showing that a stable recovery of the signal parameters require at least $\Omega(\eta^{-1})$ Fourier coefficients. A relevant concept is that of the finite rate of innovation [vetterlisignal]. This is defined in terms of the number of parameters involved in a reconstruction formula for the signal in the form
$\sum_{n\in{\mathbb{Z}}}\sum_{r=0}^{R-1}c_{r,n}\psi_{n}((x-t_{n})/T).$ (3.8)
The time points $t_{n}$ at which the signal is sampled and the coefficients $c_{r,n}$ are called the degrees of freedom in the signal. The rate of innovation is then defined as a density of these degrees of freedom in the interval over which the signal is sampled. In our formulation, we may imagine that the signal is $\hat{\mu}$ and it is sampled at the time points $\ell$. However, we have not taken the viewpoint that we need to use any reconstruction formula analogous to (3.8). Indeed, we do not even think of $\hat{\mu}$ is the signal. We think of this as the Fourier coefficients of the signal $\mu$. Therefore, the idea of innovation rate is not applicable in our setting. ∎
@@
We define a probability measure $\mu$ on $\mathbb{T}$} as a convex combination of:
@@
  • Three point-mass measures at 2,0.4,1.5-2,0.4,1.5, with weights of $60/3900,120/3900,120/3900$ respectively (anomaly).
We take 3900 samples from this measure (the number of points from each part of the measure corresponding to the numerator of the weight). The samples from the measure are visualized in Figure 3 (a) as a normalized histogram. Then, we apply $\sigma_{128}$ to get an estimation of the support, as seen in Figure 3 (b). Not only do we get an idea of the support of the measure by looking at $\sigma_{128}$, but also the amplitudes of the non-atomic components of the measure. Since we are dealing with finitely many samples, we in fact are only estimating the integral in the definition of $\sigma_{n}$ as a Monte-Carlo type summation. That is, with data $\{u_{j}\}_{j=1}^{M}$ sampled randomly from $\mu$, we estimate